# Versatile photonic frequency synthetic dimensions using a single Mach-Zehnder-interferometer-assisted device on thin-film lithium niobate

### v1

@@ Discussion
In conclusion, we design a MZI-assisted double-resonator device and experimentally demonstrate its versatility and flexibility by arbitrarily tuning the MZI with DC and RF signals. The device makes better use of the interplay of the frequency and spatial domain, and covers a vast range of possible coupling in the mentioned and parameters that can be finely modified in principle. We witness the important phenomena manifesting topology of the famous Hall ladder and Creutz ladder including the band structures modified by different synthetic magnetic fields, spin-momentum locking and Aharonov-Bohm cage effect, which paves the way for simulating more complex physical models and even predicting novel phenomena. The Aharonov-Bohm cage effect may also be utilized for engineering the number of the side bands or frequency shifting.
Therefore, our design holds significant potential in building large and high-dimensional networks as both the spatial and frequency domain can be expanded to a larger scale and a higher dimension. This makes it particularly suitable for realization on integrated optics for its striking merits such as high scalability and stability. Moreover, the TFLN is an excellent platform that not only possesses these advantages, but also has high non-linear and electric-optic coefficients, which is well-suited for our approach. However, the realization may also face challenges, for example, multiple modulation on one resonator requires long modulation area which may conflict with the requirement of loss and influence the results. Despite of this, processing on TFLN is developing rapidly recently and the performance of elements has been raising, such as less propagation loss and insertion loss and increasing the precision of the beam splitter. Hence, we believe these challenges can be addressed in the near future and our work may shed light on more interesting exploration on synthetic dimensions.
@@ Methods
Our device was fabricated on an x-cut lithium niobate on insulator wafer from NANOLN, where the 400-nm lithium niobate film and a 4.7-um $\text{SiO}_{2}$ layer are on a 525-um Si substrate in order. The resonators were patterned on a Hydrogen silsesquioxane (HSQ) film with electron-beam lithography (EBL). Subsequently, 200-nm lithium niobate was etched out via an $\text{Ar}^{+}$-based inductively coupled plasma etching (ICP-RIE) process. The residual HSQ and the redeposition were removed by a wet etching process. Two layers of ultraviolet photoresist LOR5A and S1813 were spin-coated on the device and the “bridges” for electrodes to cross the waveguides were defined by markless lithography. Then 750-nm $\text{SiO}_{2}$ was deposited and followed by a lift-off process. Subsequently, 800-nm copper electrodes were plated with a similar process. Finally, the electrodes were connected to a high-frequency printed circuit board (PCB) with bonding wires. The FSR of the resonators are approximate $\Omega=2\pi\times 8.9$ GHz. The loaded Q-factor before plating electrodes was $3.7\times 10^{5}$, while after plating the load Q-factor decreased to $1.7\times 10^{5}$ due to metal absorption. More details can be found in Supplementary information.
The RF signals were produced by three tunable microwave sources (AnaPico APUASYN20) and amplified by three amplifiers (Minicircuit ZVE-3W-183+). A distributed feedback laser is injected to the device after a polarization controller through grating couplers. The output signals are collected and analyzed by either a fast oscilloscope (Keysight DCA-X 86100D) or a Fabry-Perot cavity of which the mirror reflectivity is $98\%$ and the cavity length is approximate 1.5 millimeter.

### v2

@@ Discussion
In conclusion, we design an MZI-assisted double-resonator device and experimentally demonstrate its versatility and flexibility through arbitrary tuning of the MZI using both DC and RF signals. The device effectively leverages the interplay between the frequency and spatial domains, enabling access to a wide range of coupling configurations and parameters that can be finely adjusted in principle. We observe important phenomena in the Hall ladder and Creutz ladder models, including band structures modified by synthetic magnetic fields, spin-momentum locking, and the Aharonov-Bohm cage effect. These demonstrations pave the way for simulating more complex physical models and potentially predicting novel phenomena. Furthermore, the Aharonov-Bohm cage effect may be utilized for engineering sideband multiplicity or frequency shifts.
Therefore, our design holds significant potential in building large and high-dimensional networks as both the spatial and frequency domain can be expanded to a larger scale and a higher dimension. This makes it particularly suitable for realization on integrated optics for its striking merits such as high scalability and stability [50, 51, 52]. Aside from the aforementioned models, our configuration is potential for simulating more intriguing topological models and phenomena. The modulation on the MZI leads to symmetric cross couplings in both direction along the frequency dimension and thus the off-diagonal elements in the quasimomentum-space Hamiltonian are real. It is worth noting that, we can further break this symmetry by intentionally misaligning the frequency modes of the resonators and adjust the modulation frequencies accordingly, as proposed in the Floquet topological insulator [12]. Consequently, imaginary parts in the off-diagonal elements are introduced and models such as the SSH model can be realized. Generally speaking, our frequency-spatial hybrid configuration can simulate non-interacting (single-particle) effective Hamiltonians with some exceptions—since nonadjacent resonators may not be easily coupled, long-range coupling in the spatial direction may face challenges. Models include two particle interactions (such as in Hubbard models) are out of our present scope.
Moreover, TFLN is an excellent platform that not only offers these advantages but also features high non-linearity and electro-optic coefficients, making it well-suited for our approach. However, realization may face challenges, for instance, implementing multiple modulations on a single resonator requires a long modulation region, potentially conflicting with low-loss requirements and affecting the results. Despite this challenge, TFLN processing technology is advancing rapidly, with continuous improvements in component performance, such as reduced propagation and insertion losses, and increased beam splitter precision. Consequently, we believe these challenges can be addressed in the near future, and our work may pave the way for further exploration of synthetic dimensions.
@@ Methods
[…]
The RF signals were produced by three tunable microwave sources (AnaPico APUASYN20) synchronized by the built-in reference signals and amplified by three amplifiers (Minicircuit ZVE-3W-183+). A distributed feedback laser is injected to the device after a polarization controller through grating couplers. The output signals are collected and analyzed by either a fast oscilloscope (Keysight DCA-X 86100D) or a Fabry-Perot cavity of which the mirror reflectivity is $98\%$ and the cavity length is approximate 1.5 millimeter.